7) The zero rates for three, six, nine and twelve compounding. These rates suggest that the forwa continuous compounding. Whai need simple explain please
Answer #1

Answer :

(7). Value at interest and principal at end of 12 months = 1000000 *(1+ .095 /4) = 1023750

Present value = A * e^ - rt

= 1023750*e^ - .085

= 940,326.95

Kindly re-post remaining question separately, because I'm allowed to answer only one qusetion.